# Tight Bounds for Sparsifying Random CSPs

### v1

@@ 7.2 Coarse Classification: $k$-Plentiful
Recall in Section 6, that the largest $\mathbf{AND}_{c}$ appearing in $\mathbf{R}$ mostly governed the sparsifiability of the $r$-partite complete instance (with a final factor of $n$ being decided by more subtle properties). We now construct an analogous combinatorial property in the complete setting.
@@
7.3 Rigidity, Uncontrolled Domains, and Marginal Predicates
@@
[…]
@@
Assume that $\mathbf{S}:\operatorname{Hist}_{r}^{D}\to\mathbb{Z}_{\geq 0}$ is precisely $k$-plentiful and $h\in\mathsf{T}_{d,k}(\mathbf{S})$. We say that $\mathbf{S}_{h,E}$ is uniform if for all $g\in\operatorname{Hist}^{E}_{k}$ we have that $\mathbf{S}_{h,E}(g)$ is the same value. We say that $\mathbf{S}$ itself is marginally uniform if for every $d\in D$ and $h\in\mathsf{T}_{d,k}(\mathbf{S})$ and $E\subseteq D$ such that $h$ is $(d,E)$-rigid, we have that $\mathbf{S}_{h,E}$ is uniform.
@@
Continuing Remark 7.11, note that $\operatorname{Sym}(\mathbf{R})_{h,E}$ for $h=\operatorname{hist}(0023)$ and $E=\{0,1\}$ is not constant. Therefore, it may appear that $\operatorname{Sym}(\mathbf{R})$ is not marginally uniform. However, perhaps unintuitively, $\operatorname{Sym}(\mathbf{R})$ is indeed marginally uniform. The reason is that $\mathbf{R}$ is precisely $1$-plentiful as $2$ and $3$ appear exactly once in every $t\in\operatorname{supp}(\mathbf{R})=\{0123,0023,1123\}$. We can then compute that
@@
We can compute that each $h\in\operatorname{Sym}(\mathbf{R})$ is $(d,E)$-rigid only for $(d,E)=(2,\{2\})$ and $(d,E)=(3,\{3\})$. As such, these are the only pairs for which we need to check marginal uniformity. Note that $\operatorname{Hist}^{\{2\}}_{1}$ consists of a single element $\operatorname{hist}(2)$. Thus, the marginal predicate $\operatorname{Sym}(\mathbf{R})_{h,\{2\}}$ is automatically constant for all $h\in\operatorname{Sym}(\mathbf{R})$. Likewise, the marginal predicate $\operatorname{Sym}(\mathbf{R})_{h,\{3\}}$ is constant for all $h\in\operatorname{Sym}(\mathbf{R})$. Therefore, $\operatorname{Sym}(\mathbf{R})$ is marginally uniform.
@@
The marginal uniformity of a valued relation $\mathbf{R}:D^{r}\to\mathbb{Z}_{\geq 0}$ is different. See Section 7.6 for details.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is marginally uniform, for any $\varepsilon\in(0,1)$, there is a choice of $\ell=\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ such that an iid sampling of $\ell$ constraints from $C_{\mathrm{sym}}$ is a $(1\pm\varepsilon)$ sparsifier of $C$ with high probability.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is marginally uniform, for any $\varepsilon\in(0,1)$, every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega_{\mathbf{R}}(n^{r-k})$ constraints.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is not marginally uniform, for any $\varepsilon\in(0,1)$, there exists [^9] We warn the reader that for the first time in this paper, the constructed sparsifier is not an i.i.d. sampling of the constraints. This discrepancy is explored in Section 7.6. a $(1\pm\varepsilon)$ sparsifier of $C$ which preserves $O_{\mathbf{R}}(n^{r-k+1}/\varepsilon^{2})$ constraints.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is not marginally uniform, for any $\varepsilon\in(0,O_{\mathbf{R}}(1))$, every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega_{\mathbf{R}}(n^{r-k+1})$ constraints.
We note that Theorem 7.15 Item 2 is implied by Lemma 7.5 and Theorem 7.15 Footnote 9 is implied by Lemma 7.6. For the remaining items, we prove Theorem 7.15 Item 4 and then Theorem 7.15 Item 1 below.
@@
Note that the case $k=0$ is impossible as every predicate of arity zero is uniform (since there can only be one weight). Thus, we assume that $k\geq 1$.
@@
Let $\mathbf{S}:\operatorname{Hist}_{k}^{E}\to\mathbb{Z}_{\geq 0}$ be a non-uniform predicate. Let $C_{\operatorname{sym}}$ be the complete $k$-ary symmetric instance on $n$ variables. For some $\delta=\Omega_{\mathbf{S}}(1)$, there exist assignments $\psi,\psi^{\prime}\in E^{n}$ such that $\operatorname{Ham}(\psi,\psi^{\prime})\leq n/2$ but $\mathrm{sat}_{\mathbf{S},C_{\operatorname{sym}}}(\psi)>(1+\delta)\mathrm{sat}_{\mathbf{S},C_{\operatorname{sym}}}(\psi^{\prime}).$
@@
Since $\operatorname{Sym}(\mathbf{R})$ is not marginally uniform, there exists $d\in E\subseteq D$ and $h\in\mathsf{T}_{d,k}(\operatorname{Sym}(\mathbf{R}))$ such that $h$ is $(d,E)$ rigid, but $\operatorname{Sym}(\mathbf{R})_{h,E}$ is non-uniform. Let $\Phi$ be the set of assignments $\psi\in D^{n}$ such that $_{e}\psi=\#_{e}h$ for all $e\in D\setminus E$. Consider $x\in C$ such that $\mathbf{R}(\psi_{x})>0$. By definition of $\psi$, we must have that $\operatorname{hist}(\psi_{x})\prec_{\bar{E}}h$. However, since $h$ is $(d,E)$-rigid, we must have that $\operatorname{hist}(\psi_{x})_{e}=h_{e}$ for all $e\in D\setminus E$. Therefore, any $x\in C$ which is satisfied by $\psi$ must contain $\psi^{-1}(D\setminus E)$ which has size rkr-k. Let $C_{\operatorname{sym}}$ be the complete symmetric instance for $\operatorname{Sym}(\mathbf{R})$.
@@
Since $\mathbf{S}$ is non-uniform, we have that $p$ is not a scalar multiple of $q$. Let $\Delta_{E}:=\{a\in\mathbb{R}^{E}_{\geq 0}:\sum_{e\in E}a_{e}=1\}$. We claim there exists $a,b\in\Delta_{E}$ such that $p(a)\neq p(b)$.
@@
Next, observe that $p-\lambda q$ is a homogeneous polynomial of degree $k$, thus if we scale all inputs by some $\lambda\in\mathbb{R}$, the output will scale by $\lambda^{n}$. Thus, since $p-\lambda q$ evaluates to zero for all $a\in\mathbb{R}^{E}$ with $\sum_{e\in E}a_{e}=1$, $p-\lambda q$ always evaluates to zero. Thus, by the Schwarz-Zippel Lemma [Zip79, Sch80], $p=\lambda q$, a contradiction of that fact that $\mathbf{S}$ is non-uniform.
@@
7.5 Proof of Theorem 7.15 Item 1 (Better upper bound for uniform case)
@@
We now have a key claim which states that $(d,U_{\psi})$-rigid for all $h\in S_{\psi}$ with $h_{d}=k$. In particular, we can use the marginal uniformity framework of Section 7.3 to understand these “uncontrolled” domain elements.
@@ 7.5.1 Sandwiching
First observe that any $h\in S_{\psi}$ with $h_{d}=k$ is $(d,k)$-tight, for if there is any $h^{\prime}\prec_{\bar{d}}h$ with $h^{\prime}_{d}\geq k+1$, then $h^{\prime}\in S_{\psi}$ showing that $\psi$ is not tight. As previously shown, $h$ is $(d,E)$-rigid, so by the hypothesis of Theorem 7.15 Item 1, we have that $\mathbf{S}_{h,E}$ is uniform. In particular, this means for all $t\sim_{E}h$, we have that $\mathbf{S}(t)=\mathbf{S}(h)$. Thus, $\mathbf{S}_{E,0}(h)=\mathbf{S}_{E,1}(h)=\mathbf{S}(h)$. ∎
@@
Since $h^{\prime}$ is $(d,k)$-tight and $(d,E)$-rigid by 7.18, we can deduce that (1) $h^{\prime}\prec_{\bar{d}}g$ as $h^{\prime}(E\setminus\{d\})=0$ and (2) $\mathbf{S}_{h^{\prime},E}$ is uniform. The latter consequence implies that for all $g^{\prime}\sim_{E}h^{\prime}$ we have that $\mathbf{S}(g^{\prime})=\mathbf{S}(h^{\prime})$. Thus, we cannot have that $g\sim_{E}h^{\prime}$. In other words, there exists $e\in D\setminus E$ for which $h^{\prime}_{e}<g_{e}=h_{e}$.
Since $h^{\prime}\prec_{\bar{d}}g$, we have that $h^{\prime}_{d}-g_{d}=\sum_{e^{\prime}\in D\setminus\{d\}}g_{e^{\prime}}-h^{\prime}_{e^{\prime}}$ with the terms in the sum nonnegative. Thus, by Eq. 15, we have that $_{e}\psi=\Omega_{\mathbf{R}}(\varepsilon n).$ Since $e\not\in U_{\psi}$, there is $h^{\prime\prime}\in S_{\psi}$ with $h^{\prime\prime}_{d}=k$ and $h^{\prime\prime}_{e}\geq 1$. Thus, $h^{\prime\prime}$ is $(d,k)$-tight so $\mathbf{S}_{h^{\prime\prime},E}$ is uniform. Thus,
@@ 7.6 Independent Sparsifier Analysis
Together, the previous sections provide tight bounds on the sparsifiability of the complete CSP instance. Ultimately however, our goal is to understand the sparsifiability of a random CSP instance. In most of the cases we saw before, this distinction is without consequence, as the sparsifiers we built for the complete instance were achieved by independently, randomly sampling constraints. However, in the marginally uniform case for $\operatorname{Sym}(\mathbf{R})$, recall that we did not build our sparsifiers by independent random sampling (and instead we grouped constraints before sampling). In this section, we will identify a sub-case when i.i.d sampling suffices for sparsification, and also show that the associated condition is necessary for any non-trivial sparsification of random instances.
@@
Assume that $\mathbf{R}:D^{r}\to\mathbb{Z}_{\geq 0}$ is precisely $k$-plentiful. We say that $\mathbf{R}$ is marginally uniform if for all $d\in E\subseteq D$ and $h\in\mathsf{T}_{d,k}(\operatorname{Sym}(\mathbf{R}))$ for which $h$ is $(d,E)$-rigid, we have that for all $s\in D^{r}$ with $\operatorname{hist}(s)=h$ and all $t\sim_{E}s$, we have that $\mathbf{R}(s)=\mathbf{R}(t)$.
Note that $\mathbf{R}$ being marginally uniform implies $\operatorname{Sym}(\mathbf{R})$ is as well and so is a stronger condition. Below, we show that it is a strictly stronger condition (i.e., that there are relations $\mathbf{R}$ such that $\operatorname{Sym}(\mathbf{R})$ is marginally uniform, but $\mathbf{R}$ is not).
@@
A very important fact is that there exists $\mathbf{R}$ which are not marginally uniform, but $\operatorname{Sym}(\mathbf{R})$ is marginally uniform. For VCSPs, we can consider $D=\{0,1\}$ and $r=2$ and the following predicate $\mathbf{R}_{1}:\{0,1\}^{2}\to\mathbb{Z}_{\geq 0}$.
@@
Clearly both $\mathbf{R}_{1}$ and $\operatorname{Sym}(\mathbf{R}_{1})$ are $2$-plentiful. Further, $\operatorname{Sym}(\mathbf{R}_{1})$ is marginally uniform since it is the constant function on $\operatorname{Hist}^{\{0,1\}}_{2}$. However, $\mathbf{R}_{1}$ is not marginally uniform. To see why, consider $h\in\operatorname{Hist}^{D}_{r}$ with $h_{0}=2$ and $h_{1}=0$. It is straightforward to verify that $h$ is $(0,\{0,1\})$-rigid. However, $(0,0)\sim_{\{0,1\}}(0,1)$ but $\mathbf{R}_{1}(0,0)\neq\mathbf{R}_{1}(0,1)$.
@@
Thus, $\operatorname{Sym}(\mathbf{R}_{2})$ is marginally uniform. However, $\operatorname{hist}(0022)=h$ and $0022\sim_{\{0,1\}}1022$, but $\mathbf{R}_{2}(0022)\neq\mathbf{R}_{2}(1022)$. Thus, $\mathbf{R}_{2}$ is not marginally uniform.

### v2

@@ 7.2 Coarse Classification: $k$-Plentiful
Recall in Section 6, that the largest and $\and_{c}$ appearing in $\mathbf{R}$ mostly governed the sparsifiability of the $r$-partite complete instance (with a final factor of $n$ being decided by more subtle properties). We now construct an analogous combinatorial property in the complete setting.
@@
7.3 Rigidity, Uncontrolled Domains, and Marginal Balance
@@
Recall in Theorem 3.1, when $c=0$, we had separate behavior for $R=\{0,1\}^{r}$ and $R\neq\{0,1\}^{r}$. This distinction in the non-Boolean domain is captured by the marginal balance of $\mathbf{S}$. This definition can also be viewed as an analogue of Definition 6.3 in the uniform model.
@@
Assume that $\mathbf{S}:\operatorname{Hist}_{r}^{D}\to\mathbb{Z}_{\geq 0}$ is precisely $k$-plentiful and $h\in\mathsf{T}_{d,k}(\mathbf{S})$. We say that $\mathbf{S}_{h,E}$ is balanced if for all $g\in\operatorname{Hist}^{E}_{k}$ we have that $\mathbf{S}_{h,E}(g)$ is the same value; otherwise $\mathbf{S}_{h,E}$ is imbalanced. We say that $\mathbf{S}$ itself is marginally balanced if for every $d\in D$ and $h\in\mathsf{T}_{d,k}(\mathbf{S})$ and $E\subseteq D$ such that $h$ is $(d,E)$-rigid, we have that $\mathbf{S}_{h,E}$ is balanced.
@@
Continuing Remark 7.11, note that $\operatorname{Sym}(\mathbf{R})_{h,E}$ for $h=\operatorname{hist}(0023)$ and $E=\{0,1\}$ is not constant. Therefore, it may appear that $\operatorname{Sym}(\mathbf{R})$ is not marginally balanced. However, perhaps unintuitively, $\operatorname{Sym}(\mathbf{R})$ is indeed marginally balanced. The reason is that $\mathbf{R}$ is precisely $1$-plentiful as $2$ and $3$ appear exactly once in every $t\in\operatorname{supp}(\mathbf{R})=\{0123,0023,1123\}$. We can then compute that
@@
We can compute that each $h\in\operatorname{Sym}(\mathbf{R})$ is $(d,E)$-rigid only for $(d,E)=(2,\{2\})$ and $(d,E)=(3,\{3\})$. As such, these are the only pairs for which we need to check for marginal balance. Note that $\operatorname{Hist}^{\{2\}}_{1}$ consists of a single element $\operatorname{hist}(2)$. Thus, the marginal predicate $\operatorname{Sym}(\mathbf{R})_{h,\{2\}}$ is automatically constant for all $h\in\operatorname{Sym}(\mathbf{R})$. Likewise, the marginal predicate $\operatorname{Sym}(\mathbf{R})_{h,\{3\}}$ is constant for all $h\in\operatorname{Sym}(\mathbf{R})$. Therefore, $\operatorname{Sym}(\mathbf{R})$ is marginally balanced.
@@
The marginal balance of a valued relation $\mathbf{R}:D^{r}\to\mathbb{Z}_{\geq 0}$ is different. See Section 7.6 for details.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is marginally balanced, for any $\varepsilon\in(0,1)$, there is a choice of $\ell=\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$ such that an iid sampling of $\ell$ constraints from $C_{\mathrm{sym}}$ is a $(1\pm\varepsilon)$ sparsifier of $C$ with high probability.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is marginally balanced, for any $\varepsilon\in(0,1)$, every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega_{\mathbf{R}}(n^{r-k})$ constraints.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is not marginally balanced, for any $\varepsilon\in(0,1)$, there exists [^10] We warn the reader that for the first time in this paper, the constructed sparsifier is not an i.i.d. sampling of the constraints. This discrepancy is explored in Section 7.6. a $(1\pm\varepsilon)$ sparsifier of $C$ which preserves $O_{\mathbf{R}}(n^{r-k+1}/\varepsilon^{2})$ constraints.
@@
    If $\operatorname{Sym}(\mathbf{R})$ is not marginally balanced, for any $\varepsilon\in(0,O_{\mathbf{R}}(1))$, every $(1\pm\varepsilon)$ sparsifier of $C$ must preserve $\Omega_{\mathbf{R}}(n^{r-k+1})$ constraints.
We note that Theorem 7.15 Item 2 is implied by Lemma 7.5 and Theorem 7.15 Footnote 10 is implied by Lemma 7.6. For the remaining items, we prove Theorem 7.15 Item 4 and then Theorem 7.15 Item 1 below.
@@
Note that the case $k=0$ is impossible as every predicate of arity zero is balanced (since there can only be one weight). Thus, we assume that $k\geq 1$.
@@
Let $\mathbf{S}:\operatorname{Hist}_{k}^{E}\to\mathbb{Z}_{\geq 0}$ be an imbalanced predicate. Let $C_{\operatorname{sym}}$ be the complete $k$-ary symmetric instance on $n$ variables. For some $\delta=\Omega_{\mathbf{S}}(1)$, there exist assignments $\psi,\psi^{\prime}\in E^{n}$ such that $\operatorname{Ham}(\psi,\psi^{\prime})\leq n/2$ but $\mathrm{sat}_{\mathbf{S},C_{\operatorname{sym}}}(\psi)>(1+\delta)\mathrm{sat}_{\mathbf{S},C_{\operatorname{sym}}}(\psi^{\prime}).$
@@
Since $\operatorname{Sym}(\mathbf{R})$ is not marginally balanced, there exists $d\in E\subseteq D$ and $h\in\mathsf{T}_{d,k}(\operatorname{Sym}(\mathbf{R}))$ such that $h$ is $(d,E)$ rigid, but $\operatorname{Sym}(\mathbf{R})_{h,E}$ is imbalanced. Let $\Phi$ be the set of assignments $\psi\in D^{n}$ such that $_{e}\psi=\#_{e}h$ for all $e\in D\setminus E$. Consider $x\in C$ such that $\mathbf{R}(\psi_{x})>0$. By definition of $\psi$, we must have that $\operatorname{hist}(\psi_{x})\prec_{\bar{E}}h$. However, since $h$ is $(d,E)$-rigid, we must have that $\operatorname{hist}(\psi_{x})_{e}=h_{e}$ for all $e\in D\setminus E$. Therefore, any $x\in C$ which is satisfied by $\psi$ must contain $\psi^{-1}(D\setminus E)$ which has size rkr-k. Let $C_{\operatorname{sym}}$ be the complete symmetric instance for $\operatorname{Sym}(\mathbf{R})$.
@@
Since $\mathbf{S}$ is imbalanced, we have that $p$ is not a scalar multiple of $q$. Let $\Delta_{E}:=\{a\in\mathbb{R}^{E}_{\geq 0}:\sum_{e\in E}a_{e}=1\}$. We claim there exists $a,b\in\Delta_{E}$ such that $p(a)\neq p(b)$.
@@
Next, observe that $p-\lambda q$ is a homogeneous polynomial of degree $k$, thus if we scale all inputs by some $\lambda\in\mathbb{R}$, the output will scale by $\lambda^{n}$. Thus, since $p-\lambda q$ evaluates to zero for all $a\in\mathbb{R}^{E}$ with $\sum_{e\in E}a_{e}=1$, $p-\lambda q$ always evaluates to zero. Thus, by the Schwarz-Zippel Lemma [Zip79, Sch80], $p=\lambda q$, a contradiction of that fact that $\mathbf{S}$ is imbalanced.
@@
7.5 Proof of Theorem 7.15 Item 1 (Better upper bound for the balanced case)
@@
We now have a key claim which states that $(d,U_{\psi})$-rigid for all $h\in S_{\psi}$ with $h_{d}=k$. In particular, we can use the marginal balance framework of Section 7.3 to understand these “uncontrolled” domain elements.
@@ 7.5.1 Sandwiching
First observe that any $h\in S_{\psi}$ with $h_{d}=k$ is $(d,k)$-tight, for if there is any $h^{\prime}\prec_{\bar{d}}h$ with $h^{\prime}_{d}\geq k+1$, then $h^{\prime}\in S_{\psi}$ showing that $\psi$ is not tight. As previously shown, $h$ is $(d,E)$-rigid, so by the hypothesis of Theorem 7.15 Item 1, we have that $\mathbf{S}_{h,E}$ is balanced. In particular, this means for all $t\sim_{E}h$, we have that $\mathbf{S}(t)=\mathbf{S}(h)$. Thus, $\mathbf{S}_{E,0}(h)=\mathbf{S}_{E,1}(h)=\mathbf{S}(h)$. ∎
@@
Since $h^{\prime}$ is $(d,k)$-tight and $(d,E)$-rigid by 7.18, we can deduce that (1) $h^{\prime}\prec_{\bar{d}}g$ as $h^{\prime}(E\setminus\{d\})=0$ and (2) $\mathbf{S}_{h^{\prime},E}$ is balanced. The latter consequence implies that for all $g^{\prime}\sim_{E}h^{\prime}$ we have that $\mathbf{S}(g^{\prime})=\mathbf{S}(h^{\prime})$. Thus, we cannot have that $g\sim_{E}h^{\prime}$. In other words, there exists $e\in D\setminus E$ for which $h^{\prime}_{e}<g_{e}=h_{e}$.
Since $h^{\prime}\prec_{\bar{d}}g$, we have that $h^{\prime}_{d}-g_{d}=\sum_{e^{\prime}\in D\setminus\{d\}}g_{e^{\prime}}-h^{\prime}_{e^{\prime}}$ with the terms in the sum nonnegative. Thus, by Eq. 15, we have that $_{e}\psi=\Omega_{\mathbf{R}}(\varepsilon n).$ Since $e\not\in U_{\psi}$, there is $h^{\prime\prime}\in S_{\psi}$ with $h^{\prime\prime}_{d}=k$ and $h^{\prime\prime}_{e}\geq 1$. Thus, $h^{\prime\prime}$ is $(d,k)$-tight so $\mathbf{S}_{h^{\prime\prime},E}$ is balanced. Thus,
@@ 7.6 Independent Sparsifier Analysis
Together, the previous sections provide tight bounds on the sparsifiability of the complete CSP instance. Ultimately however, our goal is to understand the sparsifiability of a random CSP instance. In most of the cases we saw before, this distinction is without consequence, as the sparsifiers we built for the complete instance were achieved by independently, randomly sampling constraints. However, in the marginally balanced case for $\operatorname{Sym}(\mathbf{R})$, recall that we did not build our sparsifiers by independent random sampling (and instead we grouped constraints before sampling). In this section, we will identify a sub-case when i.i.d sampling suffices for sparsification, and also show that the associated condition is necessary for any non-trivial sparsification of random instances.
@@
Assume that $\mathbf{R}:D^{r}\to\mathbb{Z}_{\geq 0}$ is precisely $k$-plentiful. We say that $\mathbf{R}$ is marginally balanced if for all $d\in E\subseteq D$ and $h\in\mathsf{T}_{d,k}(\operatorname{Sym}(\mathbf{R}))$ for which $h$ is $(d,E)$-rigid, we have that for all $s\in D^{r}$ with $\operatorname{hist}(s)=h$ and all $t\sim_{E}s$, we have that $\mathbf{R}(s)=\mathbf{R}(t)$.
Note that $\mathbf{R}$ being marginally balanced implies $\operatorname{Sym}(\mathbf{R})$ is as well and so is a stronger condition. Below, we show that it is a strictly stronger condition (i.e., that there are relations $\mathbf{R}$ such that $\operatorname{Sym}(\mathbf{R})$ is marginally balanced, but $\mathbf{R}$ is not).
@@
A very important fact is that there exists $\mathbf{R}$ which are not marginally balanced, but $\operatorname{Sym}(\mathbf{R})$ is marginally balanced. For VCSPs, we can consider $D=\{0,1\}$ and $r=2$ and the following predicate $\mathbf{R}_{1}:\{0,1\}^{2}\to\mathbb{Z}_{\geq 0}$.
@@
Clearly both $\mathbf{R}_{1}$ and $\operatorname{Sym}(\mathbf{R}_{1})$ are $2$-plentiful. Further, $\operatorname{Sym}(\mathbf{R}_{1})$ is marginally balanced since it is the constant function on $\operatorname{Hist}^{\{0,1\}}_{2}$. However, $\mathbf{R}_{1}$ is not marginally balanced. To see why, consider $h\in\operatorname{Hist}^{D}_{r}$ with $h_{0}=2$ and $h_{1}=0$. It is straightforward to verify that $h$ is $(0,\{0,1\})$-rigid. However, $(0,0)\sim_{\{0,1\}}(0,1)$ but $\mathbf{R}_{1}(0,0)\neq\mathbf{R}_{1}(0,1)$.
@@
Thus, $\operatorname{Sym}(\mathbf{R}_{2})$ is marginally balanced. However, $\operatorname{hist}(0022)=h$ and $0022\sim_{\{0,1\}}1022$, but $\mathbf{R}_{2}(0022)\neq\mathbf{R}_{2}(1022)$. Thus, $\mathbf{R}_{2}$ is not marginally balanced.